# latex‘!h’ float specifier changed to ‘!ht’

### v1

@@ 3 Theoretical results for deep linear GCNs
GCNs incorporate the graph information in the layers and do not require any penalization. Instead, it is sufficient to minimize the least-squares objective over the class of GCNs,
$\mathbf{\hat{f}}_{L}:=\mathop{\rm argmin}_{g\in{\mathcal{F}}_{\text{GCN}}(L)}\big{\|}\mathbf{Y}-g(\mathbf{Y})\big{\|}_{2}^{2}=\widehat{W}P^{L}\mathbf{Y}$
@@
To understand the role of the parameter in the GCN, we disregard its randomness and focus on a fixed parameter $W.$ For any matrix $P\in{\mathbb{R}}^{n\times n},$ using that the measurement noise ${\varepsilon}_{i}$ in Model (1) is centered and the triangle inequality, we then obtain the bias-variance type decomposition
${\mathbb{E}}\left[{\|WP\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$ $={\mathbb{E}}\left[{\|WP(\mathbf{f^{*}}+{\bm{{\varepsilon}}})-\mathbf{f^{*}}\|_{2}^{2}}\right]$
$=\|(WP-I_{n})\mathbf{f^{*}}\|_{2}^{2}+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$
$\leq\Big{(}W\big{\|}(P-I_{n})\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right].$ (7)
Optimizing over the parameter $W$ balances squared bias and variance. In case of a small signal $\mathbf{f^{*}},$ the optimal $W$ will be in $[0,1)$ and induces thus shrinkage. An extreme case is $\mathbf{f^{*}}=0.$ Choosing $W=0$ will then lead to vanishing mean squared error.
In nonparametric statistics, the convergence rates of an estimator depends on the smoothness of the regression function. In the context of graphs, this smoothness is naturally expressed in terms of the neighborhood structure. As mentioned in the statistical setting, we consider a homophilic graph $\mathcal{G},$ meaning that neighboring nodes are expected to have similar signals $f_{i}^{*}\approx f_{j}^{*}$. However, the specific similarity assumption concerning neighboring nodes depends on the graph aggregation operator $P$, which incorporates various normalization approaches. We present these formally as follows.
@@
Recall that $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix, while $|\mathcal{E}|$ represents the number of edges in the graph $\mathcal{G}.$ The subsequent result bounds the mean squared error by a squared bias and a variance term.
@@
$\begin{split}\frac{1}{n}{\mathbb{E}}\left[{\|WS^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\Big{(}\underbrace{WL\Delta+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}}_{\text{bias}}\Big{)}^{2}+\underbrace{W^{2}\frac{\|S^{L}\|_{F}^{2}}{n}}_{\text{variance}},\end{split}$ (10)
@@
$\frac{1}{n}{\mathbb{E}}\left[{\|WT^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\Big{(}\underbrace{WL\Delta\sqrt{\frac{2|\mathcal{E}|}{n}+1}+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}}_{\text{bias}}\Big{)}^{2}+\underbrace{W^{2}\frac{\|T^{L}\|_{F}^{2}}{n}}_{\text{variance}}.$ (11)
While the bias term will increase in the depth $L,$ we will show that the variance term can quickly decrease in $L.$ This elucidates why, in practical applications, there is an optimal layer $L$ for implementing linear GCNs. This finding will also be further validated and demonstrated in the experimental study section (see Figure 1 in Section 4).
Besides this bias-variance trade-off phenomenon involving $L$, the variance decay can exhibit significantly different behaviors depending on the specific structure of the graph $\mathcal{G}$. Even within a single fixed graph, nodes with distinct local features may exhibit varying patterns of variance decay, which we will discuss in more detail shortly.
GCNs with linear layers denoise the response variables by averaging over neighboring nodes. A natural competitor is the estimator that estimates the $i$-th signal $f_{i}^{*}$ by uniformly averaging over the set ${\mathcal{N}}^{L}(i),$ denoting the neighboring nodes within a distance of at most $L$ from node $i$. Writing $|{\mathcal{N}}^{L}(i)|$ for its cardinality, the local average estimator is
@@
$\overline{f}_{L,i}:=\frac{1}{|{\mathcal{N}}^{L}(i)|}\sum_{j\in{\mathcal{N}}^{L}(i)}Y_{j},\quad i=1,\ldots,n.$ (13)
@@
$\frac{1}{n}{\mathbb{E}}\left[{\|\mathbf{\overline{f}}_{L}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq(\underbrace{L\Delta}_{{\text{bias}}})^{2}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|{\mathcal{N}}^{L}(i)|}}_{\text{variance}}.$ (14)
If we compare the upper bounds given by Theorems 1 and 2, we see that the bias term in (10) can be smaller depending on $W$. We analyse the variance term in details in the next session.
@@ 3.1 Bounds for the variance via weighted local walks on the graph
We now examine the dependence of the network depth $L$ on the variance decay for different local graph topologies and the graph aggregation operators $S$ and $T$.
The definition of $T$ in (3) and $S$ in (4) are based on the adjacency matrix with added self-loops $\tilde{A}.$ While the original graph is denoted by $\mathcal{G}$, we write $\mathcal{G}^{\prime}$ for the graph augmented with self-loops. Recall that the edge degrees of the graph $\mathcal{G}$ are denoted by $d_{1},\ldots,d_{n}.$ A path from node $i$ to node $j$ via nodes $\ell_{1},\ldots,\ell_{L-1}\in\{1,\ldots,n\}$ is denoted by
@@
The length of the path is $L.$ We associate to this path the weights
@@
and
$\widetilde{\omega}_{ij}(\ell_{1},\ldots,\ell_{L-1}):=\Big{(}\frac{d_{i}+1}{d_{j}+1}\Big{)}^{1/2}\omega_{i}(\ell_{1},\ldots,\ell_{L-1}).$ (15)
Let $\mathcal{P}_{L}(i\to j)$ denote all paths from node $i$ to node $j$ of length $L$ in the self-loop augmented graph $\mathcal{G}^{\prime}.$ There is a one-to-one correspondence with the paths from $i$ to $j$ of length $\leq L$ in the original graph $\mathcal{G}.$ The key observation is that for any vector $\mathbf{v}=(v_{1},\ldots,v_{n})^{\top},$
@@
see also the supplement for a derivation.
@@
The edge degrees and the length of the paths $L$ decrease the weights but increase the number of paths. Interestingly, the variance can exhibit significantly different behaviors depending on the local graph structure around node $i$. In the following, we primarily focus on the analysis for the case $P=S^{L}$. The results for $P=T^{L}$ are similar to those for $P=S^{L}$ due to the connection in (15); therefore, we have included the corresponding results and proofs in the Appendix.
Since the row sums of the matrix $S=\tilde{D}^{-1}\tilde{A}$ are all one, $S$ is a transition matrix of a Markov chain. Thus, $S^{L}$ is a transition matrix with row sums equal to one and for any node $i\in\{1,\ldots,n\},$
@@
Recall that ${\mathcal{N}}^{L}(i)$ and $|{\mathcal{N}}^{L}(i)|$ denote the set of neighboring nodes within a distance of at most $L$ from node $i$ and its cardinality. In the previous display, $\sum_{j=1}^{n}$ can be replaced by $\sum_{j\in{\mathcal{N}}^{L}(i)}.$ Applying the Cauchy-Schwarz inequality (AM-QM) and Theorem 2, we deduce
@@
For $L=1,$ equality holds in (16). The inequality shows that, in general, the local averaging estimator has smaller variance if we take the same $L$ in the GCN estimator and the local averaging estimator. On the contrary, the GCN estimator is expected to have a smaller bias, as it computes a weighted sum over the local neighborhood ${\mathcal{N}}^{L}(i),$ assigning more mass to nodes that are closer to $i$ in path distance. Although the gain in the bias is hard to quantify theoretically, we see that it outweighs the loss in the variance in our empirical studies. For instance, in all the scenarios in Figure 4 the GCN estimator has smaller test error.
We now provide an upper bound assuming that node $i$ is the root of a locally-rooted tree. In this scenario, the variance term decays exponentially in the number of layers $L$ (see Figure 2).
@@
In this case, the variance decay in $d$ is $(d+1)^{-L}.$ This is the sharp rate, since the lower bound in (16) gives the same rate, namely $1/|{\mathcal{N}}^{L}(i)|=(d-1)/(d^{L+1}-1)\approx d^{-L}.$
A slow rate of the variance term occurs if node $i$ has edge degree say $d$ but is connected to nodes with small edge degree.
@@
This indicates that under this scenario, the decay of the variance term in $d$ is significantly slower than $(d+1)^{-L}.$
Aside from the tree structure, the decay of the variance term can suddenly exhibit slow decay once an additional cycle is attached to some node $i$ in the graph $\mathcal{G}$.
@@
In practical applications, such as social network analysis, it is common to observe nodes exhibiting an additional cycle, or even multiple ones. Once a cycle is attached to some nodes, the variance decay slows down. However, increasing the number of convolutional layers may also increase the bias, as indicated by a combination of Theorem 1 and Proposition 3. Our example serves as a prototype to demonstrate why, in practice, simply stacking the layers of a GCN leads to poor performance for the resulting estimator. This phenomenon, known as over-smoothing in GCN literature [23, 3, 27, 20], contrasts with Fully Connected Neural Networks (FNNs), where theory and evidence suggest that greater depth improves representation power [31, 28]. In [23, 3], the over-smoothing phenomenon is studied asymptotically, assuming the number of layers $L$ tends to infinity. Some results in [23] also rely on large graphs with $n$ being large. Our result, established in a non-asymptotic regression setting, differs from most existing work focused on node classification. By starting with a bias-variance decomposition and analyzing weighted local walks on the graph, we quantitatively uncover the underlying reasons for this phenomenon, offering insights for improving GCNs in practice.
@@ 4 Numerical simulations
We validate our results through a set of numerical experiments. Specifically, to evaluate the link between neighborhood topology — particularly degree distribution and degree skewness — and optimal neighborhood size, we generate graphs with $n=100$ nodes using the following topologies:
@@
  • Latent variable graphs: we generate latent variables for nodes by drawing $d=2$ dimensional latent variables from a uniform distribution. The edges of the network are then sampled from a Bernoulli distribution with $p_{ij}=\text{sigmoid}(5\|U_{i}-U_{j}\|_{2})$, where the scaling factor is chosen empirically to control the edge density. To further measure the effect of the density of the network on the optimal neighborhood size, we sparsify the induced network as follows. To ensure the connectedness of the graph, we first sample a minimum spanning tree. We then delete edges that are not in the MST with probability $p$, so that a higher $p$ results in a sparser graph. The resulting sparsity of the graph is highlighted in Table 1 in the Appendix.
@@
  • Preferential attachment graphs using the Holme and Kim algorithm [13], varying their local clustering probability $p$. In this case, higher values of $p$ yield graphs with higher clustering coefficients.
@@
  • k-regular trees, where $k$ denotes the degree of each node in the tree.
@@
  • Barbell graphs, consisting of two cliques of $m$ fully connected nodes, joined by a chain on $100-2m$ nodes.
@@
In the first case, the signal $Y$ is created as a smooth function of the latent variable $U$: $Y=5\cos(U\beta)$, with $\beta$ a coefficient of the form $\beta=(\alpha,-\alpha)$, where $\alpha$ is adjusted to control different levels of smoothness across the graph: as highlighted in Figure 1 in the Appendix, lower values of $\alpha$ tend to create smooth functions, while higher values of $\alpha$ induce faster varying signals. In the three other cases, the signal $Y$ is created in a similar fashion, replacing the latent variables $U$ with the first two eigenvectors of the graph Laplacian. An example of our generation setup is provided in the Appendix.
Figure 1: Optimal number of convolutions as a function of the roughness of $\mathbf{f}^{*}$ on the latent variable graph.
Results
The results for the various topologies are provided in the Appendix. Overall, the optimal number of convolutions seems to be dependent on the local topology of the node’s neighborhood. We also assess the decay rate of the variance in rooted trees as a function of the degree. As detailed in Proposition 1, we expect the variance decay to be larger as $d$ increases. Figure 2 shows the decay of the squared bias and variance as a function of $L$ (the neighborhood depth, here logged on the x-axis) for $\mathbf{f}^{*}=5\cos(U\times\beta),$ where $\beta$ controls the smoothness of the function $\mathbf{f}^{*}$. As expected, we note that the log of the variance decays as a linear function of the number of convolutions $L$, with higher degrees yielding higher decay slopes.
Figure 2: Decay of the variance at the root of the tree as a function of $L$ (the neighborhood size) for $f^{*}=5\cos(U\beta)$, coloured by value of the degree. $\beta=(-1,1)$ induces an average smoothness of $0.14$.
Figure 3: Decay of the variance at the root of the tree as a function of $L$ (the neighborhood size) for $f^{*}=5\cos(U\beta)$, for a tree on $k=2$ nodes with, and without cycles.
@@ Results
Next, we investigate the effect of adding local cycles, as discussed in Proposition 3. To this end, we construct a binary tree rooted at node 0, and add edges between nodes at levels 2 and 3 of the tree. We report the results in Figure 3. As expected, we note an increase of the variance.

### v2

@@ 4 THEORETICAL RESULTS FOR DEEP LINEAR GCNS
Although many variants of GCNs include some form of regularization, the original design incorporates graph information within the layers and deliberately avoids penalization (Kipf and Welling,, 2017) suggesting that minimizing the least-squares objective over the class of GCNs is sufficient:
$\mathbf{\hat{f}}_{L}:=\mathop{\rm argmin}_{g\in{\mathcal{F}}_{\text{GCN}}(L)}\big{\|}\mathbf{Y}-g(\mathbf{Y})\big{\|}_{2}^{2}=\widehat{W}P^{L}\mathbf{Y}$ (7)
@@
To understand the role of the parameter in the GCN, we disregard its randomness and focus on a fixed parameter $W\in\mathbb{R}.$ For any matrix $P\in{\mathbb{R}}^{n\times n},$ using that the measurement noise ${\varepsilon}_{i}$ in Model (1) is centered and the triangle inequality, we then obtain the bias-variance type decomposition
${\mathbb{E}}\left[{\|WP\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]$
$={\mathbb{E}}\left[{\|WP(\mathbf{f^{*}}+{\bm{{\varepsilon}}})-\mathbf{f^{*}}\|_{2}^{2}}\right]$
$=\|(WP-I_{n})\mathbf{f^{*}}\|_{2}^{2}+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right]$ (8)
$\leq\Big{(}|W|\cdot\big{\|}(P-I_{n})\mathbf{f^{*}}\big{\|}_{2}+|1-W|\cdot\|\mathbf{f^{*}}\|_{2}\Big{)}^{2}$
$\quad+W^{2}{\mathbb{E}}\left[{\|P{\bm{{\varepsilon}}}\|_{2}^{2}}\right].$
Optimizing the parameter $W$ balances the squared bias and variance. In the case of a small signal $\mathbf{f^{*}},$ the optimal $W$ will be in $[0,1)$, thereby inducing shrinkage. An extreme case occurs when $\mathbf{f^{*}}={\bm{0}}.$ Choosing $W=0$ then leads to a vanishing mean squared error.
In nonparametric statistics, the convergence rate of an estimator depends on the smoothness of the regression function. In the context of graphs, this smoothness is naturally described in terms of the neighborhood structure. As mentioned, we consider graphs $\mathcal{G}$ with the property that neighboring nodes are expected to have similar signals, that is, $f_{i}^{*}\approx f_{j}^{*}$, whenever $(i,j)\in\mathcal{E}.$ However, the specific similarity assumption concerning neighboring nodes depends on the graph aggregation operator $P$, which incorporates different normalization approaches. We present these formally as follows.
@@
Recall that $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The subsequent result bounds the mean squared error as the sum of the squared bias and variance.
@@
$\frac{1}{n}{\mathbb{E}}\left[{\|WS^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\operatorname{Bias}_{S}^{2}(L,W)+\underbrace{W^{2}\frac{\|S^{L}\|_{F}^{2}}{n}}_{\text{variance}},$
@@
$\frac{1}{n}{\mathbb{E}}\left[{\|WT^{L}\mathbf{Y}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\operatorname{Bias}_{T}^{2}(L,W)+\underbrace{W^{2}\frac{\|T^{L}\|_{F}^{2}}{n}}_{\text{variance}},$
where
$\operatorname{Bias}_{S}(L,W)=|W|\cdot L\Delta+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2},$
$\operatorname{Bias}_{T}(L,W)=|W|\cdot L\Delta\sqrt{\frac{2|\mathcal{E}|}{n}+1}+\frac{|1-W|}{\sqrt{n}}\cdot\|\mathbf{f^{*}}\|_{2}.$
While the bias terms $\operatorname{Bias}_{S}(L,W)$ and $\operatorname{Bias}_{T}(L,W)$ grow with $L$, Section 4.1 demonstrates that the variance terms, specifically $\|S^{L}\|_{F}^{2}$ and $\|T^{L}\|_{F}^{2}$, can decrease rapidly as $L$ increases. Overall, Theorem 1 indicates a bias-variance type trade-off mediated by the number of layers $L.$ It implies the existence of an optimal value of $L$ that balances these two terms, providing theoretical evidence for the over-smoothing phenomenon in the regression setting. We will further validate this theoretical finding through an experimental study in Section 5, see also Figure 1.
As shown in (7), GCNs with linear layers denoise the response variables by averaging over neighboring nodes. A natural competitor is the estimator that estimates the $i$-th signal $f_{i}^{*}$ by uniformly averaging over the set ${\mathcal{N}}^{L}(i),$ which represents the neighboring nodes within a distance of at most $L$ from node $i$. Specifically, the local average estimator of $\mathbf{f^{*}}$ is given by
@@
$\overline{f}_{L,i}:=\frac{1}{|{\mathcal{N}}^{L}(i)|}\sum_{j\in{\mathcal{N}}^{L}(i)}Y_{j},\quad i=1,\ldots,n.$
The following result bounds the mean squared error of $\mathbf{\overline{f}}_{L}$ using a similar bias-variance type decomposition.
@@
$\frac{1}{n}{\mathbb{E}}\left[{\|\mathbf{\overline{f}}_{L}-\mathbf{f^{*}}\|_{2}^{2}}\right]\leq\underbrace{L^{2}\Delta^{2}}_{\text{bias}}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|{\mathcal{N}}^{L}(i)|}}_{\text{variance}}.$
By comparing the bounds in Theorem 1 and 2, under the same smoothness assumption, the bias term $\operatorname{Bias}_{S}(L,W)$ may be smaller than that of $\mathbf{\overline{f}}_{L}$, depending on the choice of $W$. We analyze the variance term in details in the next session.
@@ 4.1 Bounds for the Variance via Weighted Local Walks on the Graph
We examine in this section how the network depth $L$ and the local network topology influence the variance decay for the aggregation operators $S$ and $T$.
To this end, we introduce walk analysis as a novel approach. Recall that the definitions of $T$ in (3) and $S$ in (4) rely on the adjacency matrix with added self-loops, $\tilde{A}$. While the original graph is denoted by $\mathcal{G}$, we write $\mathcal{G}^{\prime}$ for the graph augmented with self-loops. A path from node $i$ to node $j$ via nodes $\ell_{1},\ldots,\ell_{L-1}\in\{1,\ldots,n\}$ is denoted by
@@
The length of the path is $L.$ The weights associated with this path under the aggregation operator $S$ are defined as
@@
and under $T$ as
$\widetilde{\omega}_{ij}(\ell_{1},\ldots,\ell_{L-1}):=\Big{(}\frac{d_{i}+1}{d_{j}+1}\Big{)}^{1/2}\omega_{i}(\ell_{1},\ldots,\ell_{L-1}),$
where $d_{i}$ is the edge degree of node $i$. Let $\mathcal{P}_{L}(i\to j)$ denote all paths from node $i$ to node $j$ of length $L$ in the self-loop augmented graph $\mathcal{G}^{\prime}.$ Each path of length $L$ in the augmented graph corresponds to a path of length $\leq L$ in the original graph $\mathcal{G}.$ The edge degrees and the length of the paths $L$ decrease the weights but increase the number of paths. The key observation is that for any vector $\mathbf{v}=(v_{1},\ldots,v_{n})^{\top},$
@@
with a detailed derivation provided in the appendix.
@@
If the components of $\bm{{\varepsilon}}$ are uncorrelated and have variance one, then for $P\in\{S,T\},$ we have
$\|P^{L}\|_{F}^{2}={\mathbb{E}}\left[{\|P^{L}{\bm{{\varepsilon}}}\|_{2}^{2}}\right]=\sum_{i=1}^{n}{\mathbb{E}}\left[{\big{(}P^{L}\bm{{\varepsilon}}\big{)}_{i}^{2}}\right].$
The quantity ${\mathbb{E}}[(P^{L}\bm{{\varepsilon}})_{i}^{2}]$ thus represents the contribution of node $i$ to the variance term in Theorem 1 after $L$ applications of the propagation operator $P$.
We first compare ${\mathbb{E}}[(S^{L}\bm{{\varepsilon}})_{i}^{2}]$ with the variance $\mathop{\rm Var}\nolimits(\overline{f}_{L,i})$ of the uniform average estimator $\overline{f}_{L,i}$ defined in (11). Since the row sums of the matrix $S=\tilde{D}^{-1}\tilde{A}$ are all one, $S$ is a transition matrix of a Markov chain. Thus, $S^{L}$ is a transition matrix with row sums equal to one and for any node $i\in\{1,\ldots,n\},$
@@
In fact, in the previous display, $\sum_{j=1}^{n}$ can be replaced by $\sum_{j\in{\mathcal{N}}^{L}(i)}.$ Applying the Cauchy-Schwarz inequality (AM-QM) and Theorem 2, we deduce
@@
For $L=1,$ equality holds in (12). The inequality shows that, in general, the local averaging estimator $\mathbf{\overline{f}}_{L}$ has smaller variance if we take the same $L$ in the GCN estimator and the local averaging estimator. On the contrary, the GCN estimator is expected to have a smaller bias, as it computes a weighted sum over the local neighborhood ${\mathcal{N}}^{L}(i),$ assigning more mass to nodes that are closer to $i$ in path distance. Although the gain in the bias is hard to quantify theoretically, we see that it outweighs the loss in the variance in our empirical studies. For instance, in all the scenarios in Figure 4 the GCN estimator has smaller test error.
In the remainder of this section, we study the decay rate of ${\mathbb{E}}[(S^{L}\bm{{\varepsilon}})_{i}^{2}]$ with respect to the depth $L$. Notably, the decay can vary significantly depending on the local graph structure around node $i$. Since $S$ and $T$ behave similarly, we focus on the results for ${\mathbb{E}}[(S^{L}\bm{{\varepsilon}})_{i}^{2}]$ and delay the analysis of ${\mathbb{E}}[(T^{L}\bm{{\varepsilon}})_{i}^{2}]$ to the appendix.
We first provide an upper bound assuming that node $i$ is the root of a locally-rooted tree. In this scenario, the variance term decays exponentially in the number of layers $L$.
@@
In this case, the variance decay in $d$ is $(d+1)^{-L}.$ This is the sharp rate, since the lower bound in (12) gives the same rate, namely $1/|{\mathcal{N}}^{L}(i)|=(d-1)/(d^{L+1}-1)\approx d^{-L}$ for large $d$.
A slow rate of decay for the variance occurs if node $i$ has edge degree $d$ but is connected to nodes with small edge degrees.
@@
If there exists a path from node $i$ with degree $d$ to node $j$ via nodes $\ell_{1},\ldots,\ell_{L-1}\in\{1,\ldots,n\}$ such that $d_{\ell_{1}},\ldots,d_{\ell_{L-1}}\leq 3,$ then
@@
This shows that for large $d$ and large $L$, the variance decay is necessarily much slower than the rate $(d+1)^{-L}$ obtained in Proposition 1.
The next result shows that adding a cycle to node $i$ can immediately slow variance decay.
@@
Proposition 3 serves as a prototype illustrating why stacking many GCN layers often results in poor performance. In practical applications like social network analysis, nodes frequently exhibit additional cycles. Proposition 3 suggests that in such cases, more GCN layers are needed to balance the bias-variance trade-off. However, as shown in Theorem 1, increasing the number of layers also raises the variance term, ultimately leading to the well-known over-smoothing phenomenon (Oono and Suzuki,, 2020; Cai and Wang,, 2020; Rusch et al.,, 2023; Li et al.,, 2018). This phenomenon contrasts sharply with Fully Connected Neural Networks (FNNs), where both theory and empirical evidence suggest that greater depth increases the network expressiveness (Telgarsky,, 2016; Schmidt-Hieber,, 2020).
@@ 5 NUMERICAL SIMULATIONS
We validate our results through a series of numerical experiments [^1] The code used for the experiments in the following two sections is available at the following link: https://github.com/donnate/GNN_regression. . Specifically, to evaluate the link between neighborhood topology – particularly degree distribution and degree skewness – and optimal neighborhood size, we generate graphs with $n=100$ nodes using the following topologies:
@@
  • Latent variable graphs: For each node $i$, we generate a latent variable vector $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ by sampling from a two-dimensional uniform distribution on $[0,1]^{2}$. The edges of the network are then sampled from a Bernoulli distribution with $p_{ij}=1/(1+e^{-5\|U^{\top}_{i\cdot}-U^{\top}_{j\cdot}\|_{2}})$, where the scaling factor (here, $5$) is chosen empirically to control the edge density. To further measure the effect of the edge density on the optimal neighborhood size, we sparsify the induced network as follows. We first sample a Minimum Spanning Tree (MST) as the graph’s “backbone” to ensure the sparsified graph remains connected. We then delete edges that are not in the MST with probability $p$, so that a higher $p$ results in a sparser graph.
@@
  • Preferential attachment graphs: the graphs are generated using the Holme and Kim algorithm (Holme and Kim,, 2002), varying their local clustering probability $p$. In this case, higher values of $p$ yield graphs with higher clustering coefficients.
@@
  • k-regular trees: where $k$ denotes the degree of each node in the tree.
@@
  • Barbell graphs: consisting of two cliques of $m$ fully connected nodes, joined by a chain of $n-2m$ nodes.
For the last three graph topologies, we create a latent variable $U_{i\cdot}\in\mathbb{R}^{1\times 2}$ for each node by taking the coordinates of the first two eigenvectors of the graph Laplacian. Denoting the eigenvalue decomposition of the graph Laplacian ${\bf L}=D-A$ as ${\bf L}={\bf V}\Lambda{\bf V}^{\top}$, we define $\bf U$ as the matrix formed by the first two columns of $\bf V$, i.e., ${\bf U}={\bf V}[,1:2].$
A node signal $\bf f^{*}$ is then generated as a smooth function of the latent variable vector $U_{i\cdot}\in\mathbb{R}^{1\times 2}$:
$f_{i}^{*}=2\cos(U_{i\cdot}\beta),$ (14)
@@
where $U_{i\cdot}$ denotes the $i$-th row of ${\bf U}$ and $\beta$ is a vector of coefficients of the form $\beta=(-\alpha,\alpha)^{\top}$, with $\alpha$ adjusted to control different levels of smoothness across the graph. As highlighted in Figures 7-10, lower values of $\alpha$ tend to create smoother functions, while higher values of $\alpha$ induce faster varying signals.
The left subplots of Figures 7-10 present examples of the generated graphs, with node colors indicating the true $\mathbf{f}^{*}$. The middle and right subplots provide examples of the graph’s spectral embeddings (or latent variables) ${\bf U}$, for different values of $\alpha$, showcasing signals with various levels of smoothness. Finally, the observed response vector $\mathbf{Y}=(Y_{1},\ldots,Y_{n})$ is simply generated as:
$Y_{i}=f_{i}^{*}+\varepsilon_{i}$
where $\varepsilon_{i}\sim N(0,\sigma^{2})$ is independent Gaussian noise. Unless otherwise specified, $\sigma$ is set to 1. In all subsequent experiments, consistent with the GNN literature on transductive learning (e.g., Kipf and Welling, (2017)), we partition the dataset, using 20% of the nodes for training and a separate 20% for testing. This emulates the standard GNN setting, where only a fraction of the nodes is available for training. We compare the results of the estimators obtained by (a) smoothing the signal on neighborhoods of size $L$ (where $L$ varies), using either the GCN ($T$) or GraphSAGE ($S$). These estimators have no learned parameters and solely assess the effect of neighborhood smoothing, which we compare to the local averaging estimator $\mathbf{\overline{f}}_{L}$; and (b) trained GNNs (either the GCN or GraphSAGE) with a hidden dimension of size 8 and non-linearities. The purpose of this second set of methods is to show that our results extends to trained, non-linear GNNs.
Result 1: The optimal convolution size $L$ decreases sharply as the signal’s roughness increases.
Figure 1 displays the optimal neighborhood size, averaged over 20 experiments, as a function of the graph’s roughness (simply defined as $\sqrt{[\sum_{(i,j)\in\mathcal{E}}(f_{i}^{*}-f_{j}^{*})^{2}]/|\mathcal{E}|})$ for the latent variable graph. We observe that the optimal number of convolutions decreases as a function of the graph roughness, as long as the neighborhood is informative of the underlying signal. Similar figures for the other three graph topologies (and for various levels of noise) are presented in Figures 11-13 in the appendix. The result confirms the theoretical findings presented in Theorem 1, which illustrate the bias-variance trade-off phenomenon. Interestingly, the GraphSAGE convolution $S$ results in optimal neighborhood sizes that tend to be slightly larger than those of the usual GCN convolution, particularly for very smooth signals.
Result 2: The variance appears to be generally lower for the GCN convolution ($T$) compared to GraphSAGE ($S$).
Figures 14-19 in the appendix illustrate the bias-variance trade-off across different graph topologies as the number of convolutions increases. As expected, and as predicted by (12), the local averaging estimator consistently exhibits lower variance than the GNN convolutions ($S$ and $T$). Conversely, the bias of the local averaging operator increases more substantially as a function of $L$ than that of the GNN convolutions, particularly in cases where the signal is not extremely smooth. In the Barbell graph, for instance, local averaging shows three times the bias of GNNs for $L=4$ at $\alpha=5$. GNNs with learned weights and non-linearities behave similarly for sufficiently smooth signals. Across topologies, the signal ${f_{i}^{*}}=2\cos(U_{i\cdot}\beta)$ with $\beta=(-0.1,0.1)^{\top}$ yields smooth graphs that display a clear bias-variance trade-off. Moreover, non-linear GCNs tend to exhibit lower variance than their GraphSAGE counterparts.
Result 3: The graph topology does affect the rate of variance decay.
The derived theory (Proposition 1) predicts exponential variance decay in rooted trees with respect to the number of layers $L$. Moreover, the larger the degree $d$ of the root node, the faster the variance decay. Figure 2 shows the decay of the variance as a function of $L$ (logged on the $y$-axis), for the signal $f_{i}^{*}$ defined in (14). As expected, the log of the variance decays linearly with the number of convolutions $L$, with higher degrees resulting in steeper decay slopes.
Our theory also predicts that, for instance, nodes with a high degree but connected to low-degree nodes will exhibit a lower variance decay than nodes connected to high-degree nodes. To this end, we compare the same signal in settings where either a complete graph or a star is added to a tree. In both settings, the local average estimator is expected to behave similarly. We find (Figure 20) that the variance for the GCN convolution on the star graph does not decay effectively.
@@
Next, we investigate the impact of adding local cycles, as discussed in Proposition 3. We construct a binary tree and introduce edges between nodes at levels 2 and 3. The results in Figure 3 confirm that variance decays much more slowly with cycles.